OpenStudy (just_one_last_goodbye):

10 owl bucks if you help me

OpenStudy (just_one_last_goodbye):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. Part A: Write functions to represent Cory and Roger's collections throughout the years. (4 points) Part B: How many cars does Cory have after 6 years? How many does Roger have after the same number of years? (2 points) Part C: After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically. (4 points)

OpenStudy (anonymous):

Hi again!

OpenStudy (anonymous):

So for part A, Roger's function is relatively easy to figure out

OpenStudy (just_one_last_goodbye):

Ok ^_^ i closed my question cause i know u'll help me

OpenStudy (anonymous):

So since he is constantly adding one more and he already starts with 40, we know it will be y=x+40 Where 40 is the constant (he starts with this many) x is the rate at which he adds cars (he adds one so its 1x, but thats the same as x)

OpenStudy (just_one_last_goodbye):

thats part A?

OpenStudy (anonymous):

not all, we need cory's too

OpenStudy (anonymous):

hold on, I'm just gonna figure this out, cory's is a bit weird

OpenStudy (just_one_last_goodbye):

can u give me direct answers? :/ sorry im going to get screamed at and probably get punished physically if i dont get it done

OpenStudy (anonymous):

double check the question, because cory's is a bit weird

OpenStudy (anonymous):

sorta hard to explain whats weird

OpenStudy (just_one_last_goodbye):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. Part A: Write functions to represent Cory and Roger's collections throughout the years. (4 points) Part B: How many cars does Cory have after 6 years? How many does Roger have after the same number of years? (2 points) Part C: After approximately how many years is the number of cars that Cory and Roger have the same? Justify your answer mathematically. (4 points)

OpenStudy (just_one_last_goodbye):

im copying and pasting it

OpenStudy (anonymous):

sorry nix, thats wrong

OpenStudy (anonymous):

because think about it, if x is zero, then the whole thing would be zero. But he starts out with 15 cars so it has to be something else

OpenStudy (mathstudent55):

Cory: \(y = 15(1.2)^x\)

OpenStudy (anonymous):

The question is weird because his collection increases by 20% every year. So its increasing by 20% then increase 20% of that 20%. Essentialy its something like (15x1.2)1.2 etc.

OpenStudy (anonymous):

YES thats genius

OpenStudy (anonymous):

mathstudent is right

OpenStudy (anonymous):

so A is for cory, y=15(1.2)^x and for roger, y=x+40

OpenStudy (anonymous):

For B, just plug in 6 for x

OpenStudy (anonymous):

for both equations. So for roger it would be 46, and for cory 44.78976

OpenStudy (anonymous):

but then we can't have a bit of a car.... so i guess 44?

OpenStudy (just_one_last_goodbye):

wait which is part a?

OpenStudy (anonymous):

cory, y=15(1.2)^x and for roger, y=x+40

OpenStudy (anonymous):

its asking for the functions in A

OpenStudy (just_one_last_goodbye):

ahh so those are the functions correct?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

and for C

OpenStudy (just_one_last_goodbye):

part b?

OpenStudy (anonymous):

For B just plug in x

OpenStudy (anonymous):

i mean 6 for x

OpenStudy (just_one_last_goodbye):

so I put "plug in x"

OpenStudy (mathstudent55):

Here is the explanation of Cory's function. He starts with 15 cars. His cars go up by 20% at the end of each year. After 1 year, he has \(1.2 \times 15\) At the begining of the second year, he starts with \(1.2 \times 15\) cars. At the end of year 2, he now has \(1.2 \times (1.2 \times 15) = 15 \times (1.2)^2\) At the end of year 3, he has \(1.2 \times (1.2 \times(1.2 \times 15)) = 15 \times 1.2^3\) That pattern leads to the equation: After x years, he has y cars \(y = 15(1.2)^x\)

OpenStudy (anonymous):

NONONONON0

OpenStudy (just_one_last_goodbye):

oh so put "6 for x" in part b?

OpenStudy (anonymous):

For B you plug in 6 for x yourself, and thats the answers

OpenStudy (anonymous):

we definitely know rogers, which is 46 y=6+40

OpenStudy (just_one_last_goodbye):

y = 15(1.2)^x

OpenStudy (anonymous):

44.78976

OpenStudy (anonymous):

But, idk what your instructor wants, because you can't have like 2/3 of a car

OpenStudy (just_one_last_goodbye):

idk im just gonna submit only part a

OpenStudy (just_one_last_goodbye):

thanks anyways please medal @mathstudent55

OpenStudy (anonymous):

No problem, sorry could be that big of a help

OpenStudy (just_one_last_goodbye):

it's alright

OpenStudy (nonopro):

let's name "x" the number of years. we know that 15 +15*1.20^{x} =40+x let's try after 5 years: Roger would have 45 cars and Cory 52 ( i just guessed 5, and then I calculated 15 +15*1.20^{5} ): so, no, it has to be less than 5. let's try 4: Roger will have 44, and cory 15+31=45, so that's quite good asnwer! Let's try with 3: Roger will have 43, and Cory 15+ 25=40, so 4 years are better: after approximately 4 years. (perhaps a better strategy would be to just calculate after each year starting at 1 and going up...)

OpenStudy (nonopro):

message me back if i got it right